1- You are standing on an icy surface. Your friend throws a ball with 0.4 kg at 10 m/s velocity. If your mass is 70 kg. Calculate your velocity

a- When you hold the ball.

b- When the ball hits you and bounces back 7m/s.


2- A 50kg woman stands up in a 70kg canoe 5m long. She walks from a point of 1.m from one and to a point 1.0 m from the other end. If the resistances are ignored how far is the canoe moving during this process?


Please, Expert, I need your solution to this problem. Thank you


1
Expert's answer
2020-12-25T15:21:15-0500

1.

a) "mv=(m+M)u\to u=\frac{mv}{m+M}=\frac{0.4\cdot 10}{0.4+70}=0.057(m/s)"


b) "mv=-mv'+Mu\to u=\frac{mv+mv'}{M}=\frac{0.4\cdot 10+0.4\cdot 7}{70}=0.097(m/s)"


2.

The resultant of all external forces acting on the system is equal to zero.

This mean that the position of the C.M. of the given system does not change in the process of motion.


"\overrightarrow{r}_c=const\to \Delta\overrightarrow{r}_c=0\to \sum m_i\Delta\overrightarrow{r}_i=0"


"m(\Delta\overrightarrow{r}_{mM}+\Delta\overrightarrow{r}_{M})+M\Delta\overrightarrow{r}_{M}=0"


"m(\overrightarrow{l}'+\overrightarrow{l})+M\overrightarrow{l}=0\to \overrightarrow{l}=-\frac{m\overrightarrow{l}'}{m+M}"


"l=-\frac{ml'}{m+M}=-\frac{50\cdot 3}{50+70}=-1.25(m)"